It takes Tom 4 hours to buid a fence. If he hires Jack to help him, together they can do the job in just 3 hours. If Jack built the same fence alone, how long would it take him?
Answer by stanbon(75887)

It takes Tom 4 hours to buid a fence. If he hires Jack to help him, together they can do the job in just 3 hours. If Jack built the same fence alone, how
long would it take him?
Tom Data:
time to do job= 4 hours; rate =1/4 job/hr.
Jack Data:
time to do job= x hrs ; rate = 1/x job/hr
Together Data:
time to do job= 3 hrs.; rate = 1/3 job/hr
EQUATION:
rate + rate = rate
1/4 + 1/x = 1/3
3x + 12 = 4x
x=12 hrs. (Time is would take Jack to do job alone)
Cheers,
Stan H.